OpenStudy (anonymous):

If a=x, then a+a=a+x, 2a=a=x, 2a-2x=(a+x)-2x, 2(a-x)=a-x, then 2=1?? How is this possible?

OpenStudy (anonymous):

Sorry,in the second line its 2a=a+x. My apologies.

Directrix (directrix):

If a = x, then a - x = 0. When you take the expression 2(a-x)=a-x and divide by (a -x), you are dividing by 0. Division by zero is undefined. Therefore, you cannot correctly conclude that 2 = 1.

OpenStudy (anonymous):

Thanks.